OpenStudy (anonymous):

-2x - 18 = 6(1 - 2x) + 4x x =

OpenStudy (anonymous):

put similar terms on one side what you get??

OpenStudy (kamille):

-2x - 18 = 6(1 - 2x) + 4x -2x-18=6-12+4x now you need to put all numbers with have "x" to left, and others, which don't have x to right.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes, continue